# CleAR: Robust Context-Guided Generative Lighting Estimation for Mobile Augmented Reality

### v1

@@ 4.1. Overview
We design CleAR to tackle the abovementioned challenges and provide robust and high-quality HDR environment maps for mobile AR. The design of CleAR leverages the recent advancement of generative models (Zhang et al., 2023), which can better integrate with rich conditioning inputs to ensure appropriate image context on the generated environment maps. In contrast to prior works that focus on generative model architectures (Phongthawee et al., 2023; Wang et al., 2022; Akimoto et al., 2022; Gardner et al., 2017), the core of CleAR lies in the novel integration between image generative models and AR systems to ensure high-quality and robust generative lighting estimation. Specifically, CleAR uses a context-guided generative lighting estimation pipeline (§4.2) that incorporates mobile AR context information to estimate visually coherent environment maps. Additionally, CleAR uses an edge-device collaborative estimation architecture (§4.3) to allow robust and temporally consistent lighting estimation. On the edge, we leverage a multi-output estimation strategy (§4.3.1) to tackle the inherent ambiguity of generative models in environment lighting estimation. On the client, we design several techniques to effectively select and refine (§4.3.2) edge-estimated environment maps to match real-time environment lighting conditions observed on the AR device.
@@ 4.2. Context-Guided Generative Lighting Estimation
Estimating HDR omnidirectional lighting from limited LDR environment observations is a process that requires significantly increased information. While generative models are good at such tasks, properly controlling the generation process is critical to ensure accurate results when generating environmental lighting. Our key insight to address this issue is leveraging AR context data to control the generation. Figure 3 shows an overview of our context-guided generative lighting estimation pipeline.
Figure 3. Overview of CleAR. CleAR uses several types of AR context data to guide a two-step high-quality lighting estimation pipeline. The first step (§4.2.1) completes partial environment observations in the LDR pixel range domain. The completed environment maps will be post-processed and selected with our adaptive estimation control and refinement components (§4.3) and given to the following high-intensity map estimation step (§4.2.2). Finally, CleAR outputs an HDR environment map by combining the completed LDR environment map and the high-intensity map.
@@
CleAR completes the environment map from limited camera observations in the LDR image pixel ranges at the first step. Here, we follow the standard image format convention and consider the pixel range between $[0,255]$ as the LDR pixel range. To achieve the LDR completion, we leverage ControlNet (Zhang et al., 2023) and pre-trained StableDiffusion (Rombach et al., 2021) models to train controllable image generative models. We chose ControlNet because it provides flexible controllability in image generation with text and image conditioning data that can be naturally integrated with the AR context. The pre-trained StableDiffusion model provides embedded image generation knowledge learned from large-scale image generation datasets  [^1] Note that CleAR’s design is not tied to ControlNet nor StableDiffusion, and that means other conditional image generative models can serve as drop-in replacements to improve the performance of CleAR in the future.. While the pre-trained large models can natively generate diverse images, they must be fine-tuned to output the correct environment map content and styles. To fine-tune the completion model, we generated approximately 30K data items from online sources. The data collection and generation processes will be described in §5.
In the completion process, we provide the ControlNet with camera RGB observations and scene semantics as image conditioning inputs and encode ambient light intensity and color temperatures in the text conditioning input. For camera RGB observations, we stitch multi-view camera RGBs into 360 panoramic environment map images in advance. Note that this design can be extended with more sophisticated environment reconstruction methods like (Zhao et al., 2022) or combined with multi-user environment observation sharing to increase the observed environments. The environment semantic map is also processed into 360 panoramic images as a ControlNet input. We follow the ADE20K (Zhou et al., 2017) definition and use 150 instance labels in the semantic map data.
For encoding ambient lighting condition data, we convert numerical lighting condition data into words describing the lighting condition’s characteristics. We use the following text prompt template to describe ambient lighting conditions. We create the ambient light property labels based on partial environment map image pixels with empirically derived threshold values. The light intensity label is defined by the mean pixel intensity calculated from Equation (1) with values between 0.25 and 0.40 as neutral, and the rest values as dark and bright. The color temperature label is defined by the mean pixel color temperature with values between 3500K and 5500K as neutral, and the rest values as warm and cool (Noguchi and Sakaguchi, 1999). The lighting condition words are marked in bold font.
@@
In Figure 4, we show examples of environment map high-intensity components (high-intensity map). The high-intensity components describe environment light intensities and directionalities, which are critical for rendering photorealistic highlight and shadow effects on AR objects. However, due to the device’s camera imaging capability, this information often cannot be directly captured by mobile AR devices. Thus, it is important to estimate high-intensity environment components from LDR environment observations.
Similar to the completion of the LDR environment map, we utilize an image generation model to generate high-intensity maps based on the LDR completion results. The key challenge to building a high-intensity map estimation model is data scarcity. Specifically, existing high-quality HDR environment map data are very limited. For example, the Laval dataset only includes approximately 2K data items. Training generative models on small datasets can be highly unreliable. Additionally, pre-trained backbone models, like StableDiffusion (Rombach et al., 2021), cannot be directly leveraged because they are trained on LDR images.
To address these challenges, we propose a novel learning formulation to allow high-intensity estimation in the LDR pixel range domain. Specifically, our design includes two key steps. First, we use a scaling transformation to convert HDR environment maps into LDR and high-intensity maps, like the examples in Figure 4. We define the high-intensity map as a 360 image that describes the high-intensity light positions, directions, and high-intensity pixel values. In particular, the high-intensity map is also an LDR image where the pixels describe the intensity exponent of the original HDR image pixels. we use Equation (2) to implement the conversion.
(2) $I_{m}={2.0}/({1+e^{-I_{i}}})-1.0$
Then, we fine-tune a pre-trained generative model to learn the generative process of generating high-intensity maps from LDR environment maps. We can use pre-trained generative models for this task because the input and output image pixels are in the LDR range. More importantly, the high-intensity values are strongly associated with simple LDR image features, such as bright spots and light strips, making the high-intensity map estimation task relatively simple compared to the LDR completion task. This novel design allows us to overcome the HDR data availability challenge and use the small amount of ground-truth HDR data from the Laval dataset to train the high-intensity estimation model. During training and inference, we also supply the following text prompt to provide quality conditions to the high-intensity estimation process:
@@
The prompt guides the high-intensity map estimation model in generating the correct image style, outputting consistency in high-intensity estimation, and helping to reduce visual artifacts in the estimated results. Combined with the completed LDR environment map from the first step (§4.2.1), our generative lighting estimation pipeline outputs a 360 HDR environment map.
4.3. Edge-device Collaborative Estimation
@@
Figure 5. Estimation refinement workflow. CleAR uses a hybrid architecture to offload heavy generative model inference to the edge server while adaptively refining estimation results on AR devices in real time. We adopt a multi-output estimation strategy to tackle the estimation ambiguity (§4.3.1). On the client, our system first matches the color appearances between the completed environment maps to the real-time camera observations (§4.3.2). This operation helps improve estimation accuracy in challenging lighting conditions and enables real-time adaption of the generative estimation result. Here, we show the refinement process for an extremely cool environment lighting temperature condition to highlight the importance of color appearance matching. Then, our system client chooses the best estimation result based on real-time environment observation. The chosen environment map is combined with its corresponding high-intensity map estimation to create the final HDR environment map for AR virtual object rendering.
Robust mobile AR lighting estimation requires accurate and timely estimation result updates to address the changing environmental lighting. However, providing real-time generative estimation on AR devices is impractical due to the high inference latency of generative models. Therefore, we adopt an edge-device collaborative estimation architecture design similar to recent works (Ben Ali et al., 2022; Zhao and Guo, 2021). Figure 5 illustrates the collaborative estimation process. Specifically, we utilize a multi-output estimation strategy (§4.3.1) to tackle the estimation ambiguity problem. And, on the client, CleAR selects and refines the best environment map on the AR device based on real-time environment observations (§4.3.2).
4.3.1. Multi-output Estimation Strategy.
Completing 360 environment maps from low-FoV environment observations is inherently ambiguous. To tackle this problem, we utilize a multi-output estimation strategy. The key insight of this strategy is to generate multiple completion variants and match the best one to real-time environment observations. The process can be described in three steps. First, CleAR uses different random seeds on the edge to adjust LDR completion ControlNet to output different completion results. After the CleAR client receives the completion results, the client first refines the completed environment map colors (§4.3.2). Finally, it utilizes another step on the client to match the best completion result to real-time environment observations. Specifically, our matching technique first creates color palettes from the observation and estimation environment map images. The color palettes are selected via the K-means algorithm by finding the five most common colors for each environment map image. The selected colors are profiles of each environment map. Then, our method calculates a cosine similarity between the observed and estimated environment map palettes and finds the best environment map for the current environment lighting. The total cosine similarity is calculated as the sum of per-color cosine similarity between color palettes.
4.3.2. On-device Real-time Refinement
Addressing unpredictable changing environment lighting is crucial for supporting temporally consistent AR object rendering. To address this challenge, CleAR uses a lightweight on-device refinement technique to refine the multi-output LDR completion results to align with the real-time environment observations. Specifically, our technique seeks to create a color refinement matrix of pixel color multipliers for the completed LDR environment maps. The color refinement matrix consists of two types of multiplier values: (i) the global color multiplier and (ii) the local color multiplier. The global color multiplier adjusts the overall image colors of completed environment map images, while the local color multiplier adjusts fine-grained colors on the observed environment map regions.
Our technique calculates the global color multiplier by deriving per-channel multipliers as the ratio between average colors in the estimated and observed environment map images. The local multipliers are calculated by first splitting the estimated environment map images and the partial environment observation image into $NxM$ patches. Then, our technique calculates local multipliers as the average color ratios for each patch. Finally, the two color refinement terms are combined into a color refinement matrix. The calculated local multipliers for environment map regions with observed environment pixels and the global multiplier for others. A 3x3 median blur filter is used to address color smoothness between the observation edges. Empirically, we found 8x8 is the best image patch size. As shown in Figure 6, our refinement technique can adjust environment map colors even in very challenging lighting conditions.
@@ 5.1. System Implementation
We implement CleAR based on a recent AR data streaming framework, ARFlow (Zhao and Guo, 2024), which enables low-latency bidirectional data streaming between our edge service and AR application client. CleAR consists of a client-side and a server-side component. Figure 7 shows an overview of the CleAR architecture. We implement the client as a Unity3D [^2] Unity3D: https://unity.com package. The client is responsible for streaming and receiving mobile AR device data and providing AR applications with estimated environment maps. We use ARFlow’s built-in data collection feature to capture camera RGB, depth, and device tracking data provided by ARFoundation [^3] ARFoundation: https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.0/, Unity’s low-level AR framework. The captured data is then streamed to our server via the ARFlow gRPC service. The server component is a Python-based service for the generative lighting estimation pipeline host and inference. On the server, we first decode and prepare the AR data into NumPy arrays. During generative model inference, we combine the RGB and semantics-guided ControlNet models using the MultiControlNet pipeline from the HuggingFace Diffusers [^4] Diffusers: https://huggingface.co/docs/diffusers library. We configure the ControlNet model inference to use the UniPC (Zhao et al., 2024) sampler with 20 steps. To accelerate inference speed, We include a specialized diffusion model optimization library xformers555xFormers: https://facebookresearch.github.io/xformers/ to reduce the estimation latency. Our generative estimation pipeline outputs environment maps at 512x256 resolution.
a System Architecture Overview
b AR Device View
c Edge Server Data
Figure 7. System implementation and AR application integration. 7a An overview of the software components of CleAR. 7b An AR application screenshot of a rendered virtual bunny with CleAR estimated environment lighting. 7c The device uploaded AR data to the edge server.
5.2. Dataset Generation
5.2.1. LDR Environment Map Completion Data
Next, we mask the LDR environment map images at random angles to generate AR camera observation images. The masks are generated using the pin-hole camera model to simulate the real-world partial environment observations. We also combine multiple image masks to simulate multi-view environment observations. Specifically, the number of views is randomly chosen from 1 to 5, and the camera horizontal FoV is randomly chosen between 60 and 120 degrees. For environment semantics, we use a pre-trained semantic map estimation model (Xiao et al., 2018) to estimate semantic maps directly from the collected LDR environment maps. We also use the LDR image mask to generate masked semantic maps representing environment semantic information received by AR devices. For the ambient lighting condition prompt generation, we create the ambient light property labels using the masked partial environment map image pixel values and our defined system threshold values in §4.2.1.
5.2.2. High-intensity Map Estimation Data
With our novel high-intensity map learning objective design, the high-intensity map estimation model can be fine-tuned with a small training data set. Therefore, we only use the Laval dataset’s training split to generate the high-intensity map estimation training data. In total, we created 1489 data pairs of LDR environment map images and high-intensity map images. To generate each data item, use Equation 2 to generate the data items. Specifically, we first clamp the raw HDR environment map into $[1,+\infty]$ and then scale the results to $[0,+\infty]$ to get a raw high-intensity map $I_{i}$. Next, we convert the range of the raw high-intensity $[0,+\infty]$ to the standard LDR pixel range in $[0,1]$. Note that although this transformation is not lossless, our evaluation suggests minimal impact on the overall lighting estimation. In addition to the generated image pair data, we add the text prompt P2 during training.
5.3. Model Training

### v2

@@ 4. CleAR Design
Figure 3. Overview of CleAR. CleAR uses four types of AR context data to guide a two-step high-quality lighting generation pipeline. The first step (§4.2.1) completes partial environment observations in the LDR pixel range domain. The completed environment maps will be post-processed and selected using our real-time estimation refinement components (§4.3) and then passed on to the subsequent high-intensity map estimation step (§4.2.2). Finally, CleAR outputs an HDR environment map by combining the completed LDR environment map and the high-intensity map.
@@ 4.1. Overview
We design CleAR to support robust and high-quality lighting estimation for mobile AR. CleAR estimates omnidirectional environment lighting, a 360 HDR environment map image, from multimodal AR context data. In this work, we consider four types of AR data, including low-FoV LDR images, ambient light intensity and color temperature, and scene semantic map.
At a high level, estimating HDR omnidirectional lighting from low-FoV LDR environment observations is a process of significant information increase. To tackle this challenging problem, CleAR leverages the recent advancements in conditional image generative models (Zhang et al., 2023) to create missing environment lighting information, guided by AR context data. The key design of CleAR is a novel integration of image generative models to lighting estimation systems to ensure high-quality and robust estimation results.
As illustrated in Figure 3, the design of CleAR revolves around an AR context-guided generative lighting estimation pipeline (§4.2). While generative models excel at image generation in general, producing accurate HDR environment maps requires precise conditioning to ensure visual consistency with real-world scenes, especially under complex lighting conditions and limited environmental observations. We address this challenge by using AR context data to guide generation, offering both coarse- and fine-grained conditioning.
Another critical aspect of robust mobile AR lighting estimation is the timely update of estimation results to address changes in environmental lighting. However, the high inference latency of generative models limits the frequency of lighting updates. To address this challenge, CleAR employs an estimation refinement workflow (§4.3) that adapts the color profile of estimated environment maps to real-time AR camera images, ensuring temporally consistent lighting estimation. Specifically, CleAR first employs a multi-output generation strategy (§LABEL:subsubsec:multi_output_generation) to create reasonable yet diverse estimation results for unobserved environments. Then, CleAR uses several techniques (§4.3.1 and §4.3.2) to refine the previously estimated environment maps and select the best one as the estimation output. This workflow is well-suited for the edge-assisted AR setup, where computationally intensive generative models run on edge servers, while AR devices handle lightweight, real-time refinement.
@@ 4.2. Context-Guided Generative Lighting Estimation
In the first step, CleAR completes the environment map from limited camera observations, treating the standard 8-bit pixel range as the LDR range. Our LDR completion model builds on the ControlNet (Zhang et al., 2023) architecture with pre-trained Stable Diffusion (Rombach et al., 2021) model weights. We choose ControlNet for its flexible controllability using text and image conditioning, which integrates well with the AR context. The pre-trained StableDiffusion model provides strong prior knowledge learned from large-scale image generation datasets [^1] Note that CleAR’s design is not tied to ControlNet nor Stable Diffusion; other conditional image generative models can serve as drop-in replacements to improve the performance of CleAR in the future.. These priors enable us to train environment map generation models without starting from scratch entirely, also allowing for better generalization to diverse real-world scenarios. While Stable Diffusion excels at general image generation, it must be fine-tuned on environment map data to adapt it to our task of generating panoramas in the correct format and style.
During the completion process, we encode four types of AR context data into three input modalities for ControlNet: (i) a camera RGB image as direct environmental observation, (ii) a scene semantic map for structural conditioning, and (iii) a text prompt encoding ambient light intensity and color temperature. As part of input preparation, we stitch multi-view camera RGBs into 360 panoramic environment map images in advance. This process can be extended with more sophisticated reconstruction methods like (Zhao et al., 2022a) or combined with multi-user observation sharing to increase the observed environments. Similarly, environment semantic maps are also processed into 360 panoramic images, following the ADE20K (Zhou et al., 2017) definition with 150 instance labels.
To encode ambient lighting conditions, we convert numerical sensor data into natural language descriptions of the environment’s lighting characteristics. We use the following text prompt template to describe ambient lighting conditions. The lighting condition words are marked in bold font.
@@
The ambient light property labels are created based on partial LDR image pixels with empirically derived threshold values. The light intensity label is defined by the mean pixel intensity (Bolduc et al., 2023) with values between 0.25 and 0.40 as neutral, and the rest of the values as dark and bright. We define the color temperature label based on the mean pixel color temperature: values between 3500K and 5500K are considered neutral, while values outside this range are categorized as warm or cool (Noguchi and Sakaguchi, 1999).
@@
The second step of our pipeline is to estimate high-intensity components in the completed LDR environment maps. The high-intensity components describe environment light intensities and directionalities, which are critical for rendering photorealistic highlight and shadow effects on AR objects. However, due to limited camera capabilities, mobile AR devices often fail to capture high-intensity lighting information. As shown in Figure 4, we decompose each HDR environment map into two parts using Equation (1)—an LDR representation and a high-intensity component (in the form of a high-intensity map)—for visualization. We define a high-intensity map as a 360 LDR image that describes the high-intensity light positions, directions, and pixel values from the original HDR environment map.
(1) $I_{m}={2.0}/({1+e^{-I_{i}}})-1.0$
@@
To generate high-intensity maps, we again condition a generation model on the completed LDR environment map produced in §4.2.1. One of the key challenges in building reliable high-intensity map estimation models is the scarcity of high-quality HDR environment maps. For example, the Laval dataset includes only about 2K data items. Training generative models on small datasets can be highly unreliable. Additionally, pre-trained backbone models, like Stable Diffusion (Rombach et al., 2021), cannot be directly used for high-intensity map estimation because they are trained on LDR images.
To address these challenges, we propose a novel learning formulation to allow high-intensity estimation in the LDR pixel range domain. Specifically, our design includes two key steps. First, we use a scaling transformation (Equation(1)) to convert HDR environment maps into LDR and high-intensity maps. Then, we fine-tune a pre-trained generative model, using the transformed Laval dataset, to learn the mapping from LDR environment maps to high-intensity maps. This is feasible because both input and output are represented in the LDR pixel range. More importantly, the high-intensity values often align with simple LDR image features, such as bright spots and light strips, making high-intensity map estimation a simpler task compared to the LDR completion task. During training and inference, we use the following text prompt to guide the high-intensity estimation process in producing consistent image style and reducing visual artifacts.
@@
This novel design allows us to overcome the HDR data availability challenge and use the small amount of ground-truth HDR data from the Laval dataset to train the high-intensity estimation model. Combined with the completed LDR environment map from the first step (§4.2.1), our generative lighting estimation pipeline outputs a 360 HDR environment map.
4.3. Real-time Estimation Refinement
Completing 360 environment maps from low-FoV environment observations is an inherently ambiguous process. Although AR context data helps condition the generation process, each generation can take several seconds—leading to potential mismatches between the current AR camera image and the one used during generation. To address this, we propose a multi-output estimation process that enables real-time refinement. Our key insight is to generate multiple completion variants and select the one that best matches the real-time AR view. At a high level, the process consists of three steps as illustrated in Figure 5. First, CleAR uses different random seeds in the LDR completion network to output different completion results. Next, CleAR adjusts the color appearance of the LDR environment maps based on real-time AR camera image (§4.3.1). Finally, CleAR selects the optimal result that matches the color palette of the current AR view (§4.3.2).
@@ 4.3. Real-time Estimation Refinement
Figure 5. Estimation refinement workflow. CleAR uses a hybrid architecture to offload heavy generative model inference to the edge server while adaptively refining estimation results on AR devices in real time. We adopt a multi-output generation strategy to tackle the estimation ambiguity. On the client, our system first matches the color appearances between the completed environment maps and the real-time camera observations (§4.3.1). This operation helps improve estimation accuracy in challenging lighting conditions and enables real-time adaptation of the generative estimation result. Then, our system client chooses the best estimation result based on real-time environment observation. The chosen environment map is combined with its corresponding high-intensity map estimation to create the final HDR environment map.
Our multi-output generation design not only enables CleAR to output accurate lighting estimation results, but also provides practical support for responsive lighting estimation in the real-time AR application workflow. Specifically, generative models, particularly diffusion-based models, typically suffer from long inference latency, making real-time estimation (¿30 FPS) impractical on mobile AR devices. Moreover, the high memory footprint of such models, often exceeding 2–4 GB during inference (Rombach et al., 2021), makes them unable to run on resource-constrained mobile platforms that must concurrently handle rendering and sensing tasks. In contrast, our refinement method enables the estimated environment maps to adapt to minor changes in environmental lighting, thereby reducing the need for frequent inference of the generative model. To facilitate real-time deployments, CleAR adopts an edge-device collaborative estimation architecture, similar to recent works (Ben Ali et al., 2022; Zhao and Guo, 2021). Specifically, LDR environment map completion and high-intensity map estimation can be offloaded to an edge server, while lightweight color refinement and result selection algorithms are executed on-device.
4.3.1. Color Appearance Adaptation
Addressing unpredictable changes in environmental lighting is crucial for supporting temporally consistent rendering of AR objects. Therefore, CleAR employs a lightweight on-device refinement technique to refine the multi-output LDR completion results to adapt their color appearances to real-time AR camera images. Our design selects the AR camera image as the adaptation target because the ultimate goal of mobile AR lighting estimation is to support the visually coherent insertion of AR objects into camera images. Additionally, our on-device real-time color appearance adaptation design is well-positioned to address the real-time estimation needs in mobile AR. Specifically, by applying color adaptation to the estimated environment maps, CleAR achieves consistently high estimation accuracy with few generative model inference requests. A recent work, DiffusionLight (Phongthawee et al., 2023), uses a similar strategy that has to invoke large quantities of inferences to achieve desirable quality results. However, as we will show in §6, DiffusionLight incurs a very high end-to-end inference time (359.9s), making its multi-output estimation strategy infeasible for real-time AR applications. In comparison, using our color adaptation technique, CleAR can achieve better estimation accuracy than DiffusionLight while using approximately 110X less time. Additional implementation details and testing examples of color refinement are included in Appendix §D.
4.3.2. Estimation Output Selection.
As the final step of the real-time estimation refinement process, CleAR selects the best environment map image from the previously refined LDR environment map images. In this process, CleAR searches for the best estimation candidate by identifying the highest cosine similarity between the color palettes of the observation and estimation environment map images. Specifically, the color palettes are selected using the K-means algorithm, which identifies the five most common colors for each environment map image. The total cosine similarity is calculated as the sum of per-color cosine similarity between color palettes. The selected LDR environment map will later be combined with its corresponding high-intensity map to form a complete HDR environment map for mobile AR applications.
@@ 5. Implementations
a System Architecture Overview
b Example AR Application
c Edge Server Data Visualization
Figure 6. System implementation and AR application integration. Figure 6a shows an overview of the software components of CleAR. Figure 6b depicts an AR application screenshot of a rendered virtual bunny with CleAR estimated environment lighting. Figure 6c illustrates the visualization component at the edge server for the collected AR data.
@@ 5.1. System Implementation
We implement CleAR based on a recent AR data streaming framework, ARFlow (Zhao and Guo, 2024), which enables low-latency bidirectional data streaming between the AR device and the edge server. We also implement an example object placement AR application with CleAR to demonstrate the end-to-end rendering results. Figure 6 shows an overview of the CleAR architecture. We implement the client as a Unity3D [^2] Unity3D: https://unity.com package. The client is responsible for streaming and receiving data from mobile AR devices and providing AR applications with estimated environment maps. We use ARFlow’s built-in data collection feature to capture camera RGB, depth, and device tracking data provided by ARFoundation [^3] ARFoundation: https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.0/, Unity’s low-level AR framework. The captured data is streamed to our server via the ARFlow gRPC service. The server component is a Python-based service for the generative lighting estimation pipeline. During generative model inference, we combine the RGB and other context information using the MultiControlNet pipeline from the HuggingFace Diffusers [^4] Diffusers: https://huggingface.co/docs/diffusers library to run different ControlNet models. We configure the ControlNet model inference to use the UniPC (Zhao et al., 2024) sampler with 20 steps. Our ControlNet models output environment maps at a resolution of 512x256. xformers555xFormers: https://facebookresearch.github.io/xformers/ is used to accelerate model inference.
5.2. Training and Testing Data Generation
We curate a large-scale dataset for training and testing CleAR. For LDR completion model training, we assemble 30K LDR environment map images from two prominent LDR indoor environment map sources: the Matterport3D dataset (Chang et al., 2017) and Structured3D datasets (Zheng et al., 2020). Matterport3D dataset provides real-world captured environment maps, and Structured3D dataset consists of photorealistic synthetic ones. For high-intensity map estimation model training, we use the training split of the Laval dataset (Gardner et al., 2017) to create 1,489 pairs of LDR environment map images and corresponding high-intensity map images. To evaluate CleAR under diverse lighting conditions, we create variants of the Laval test set that preserve the original environment map visual context. We generate these variants by uniformly scaling the environment map images. For light-intensity editing, all color channels are scaled equally; for color temperature editing, only the red and blue channels are adjusted (Afifi and Brown, 2020). See Appendix §B for additional details on the data generation process, and Appendix §C for details of model training.